# An Optimal Cascade Feature-Level Spatiotemporal Fusion Strategy for Anomaly Detection in CAN Bus

### v1

@@ I Introduction
  • Providing detailed results, performing statistical validation through paired t-tests and ROC curve to demonstrate the reliability and stability of the findings of the current study.
@@ II Related Works
TABLE I: Summary of Related Work
@@ III-A Datasets
TABLE II: Description of the First Dataset [12]
@@
TABLE III: Description of the Second Dataset [10]
@@ III-B Spatial Concept
TABLE VI: Feature Values Observed in CAN ID 0x490 During a Spoofing Attack [12]
@@ III-C Temporal Concept
Figure 3: Time interval analysis for CAN ID 0x490 during an attack [12].
@@ IV-B Spatial Concept
Raw Spatial Features 0.9682 0.7515 0.9080 0.8224
1D-CNN Spatial Features 0.9546 0.7211 0.7220 0.7215
Raw and 1D-CNN Spatial Features 0.9814 0.8790 0.8951 0.88705
xLSTM Spatial Features 0.9450 0.6616 0.6670 0.6643
Raw and xLSTM Spatial Features 0.9811 0.8789 0.8908 0.8848
Simple RNN Spatial Features 0.9412 0.6376 0.6468 0.6421
Raw and Simple RNN Spatial Features 0.9805 0.8752 0.8879 0.8815
GRU Spatial Features 0.9418 0.6418 0.6488 0.6453
Raw and GRU Spatial Features 0.9808 0.8783 0.8885 0.8834
LSTM Spatial Features 0.9432 0.6506 0.6571 0.6538
Raw and LSTM Spatial Features 0.9812 0.8804 0.8916 0.8859
@@
However, combining the module-generated features with raw spatial features leads to a significant enhancement in the F1-score by approximately 6% and in accuracy by around 1.5% across all models. This highlights not only the necessity of cascade feature-level fusion but also demonstrates its effectiveness. Additionally, cascade feature-level fusion enriched by xLSTM, LSTM, and 1D CNN, achieves accuracy higher than 0.9810, displaying the capability of these three models in this task.
@@ IV-C Temporal Concept
In the current subsection, the performance of raw temporal features, including Timestamp and CAN ID, as well as extracted temporal features, including Ratio and SE, and the combination of these four features are evaluated. Table IX depicts that the accuracy of the model using only raw temporal features is approximately 96%. In contrast, the extracted temporal features with a filter size of 7500 (a random choice) demonstrate around 0.3% higher accuracy. However, the combination of extracted temporal features with raw features does not show promising results, as expected. This is because the performance of the extracted features depends on the filter size, and there is a need to fine-tune this filter size. Therefore, the filter size is adjusted in the following subsection.
@@
Timestamp, CAN ID 0.9632 0.8716 0.6440 0.7407
Ratio, SE 0.9659 0.9190 0.6393 0.7540
Timestamp, CAN ID, 0.9630 0.86745 0.6467 0.7410
@@
9332 Timestamp, CAN ID, Data3, Data4, 0.9492
@@ IV-D Optimal Cascade Feature-Level Fusion
All Raw Features 0.9904 0.9342 0.9485 0.9413
1D-CNN-Based Optimal Cascade Feature-Level Fusion 0.9936 0.9620 0.9584 0.9602
xLSTM-Based Optimal Cascade Feature-Level Fusion 0.9936 0.9620 0.9578 0.9599
Simple RNN-Based Optimal Cascade Feature-Level Fusion 0.9934 0.9610 0.9581 0.9595
GRU-Based Optimal Cascade Feature-Level Fusion 0.9934 0.9614 0.9583 0.9599
LSTM-Based Optimal Cascade Feature-Level Fusion 0.9935 0.9618 0.9578 0.9597
@@ V-A Robustness and Reliability
This subsection examines the statistical evaluation of the model and analysis of its error rate, which are crucial for ensuring the reliability of the results. Table XIII presents the results of the 5x2cv paired t-test conducted to evaluate the statistical significance of the proposed model compared with the raw data. The t-statistic and p-value are equal to 20.51 and 5.09E-6, respectively. These results indicate that the proposed methodology is statistically significant. Additionally, Fig. 5 illustrates the comparison between the receiver operating characteristic (ROC) curves of the proposed model and raw data. The results show that raw features achieved an area-under-the-curve (AUC) of 0.9713, while the optimal feature space achieved an AUC of 0.9783, demonstrating that the proposed methodology outperforms in terms of error rate.
TABLE XIII: Statistical Analysis of Proposed Optimal Cascade Feature-Level Fusion Method Using the 5x2cv Paired T-Test.
@@
Figure 5: Comparison of ROC curves: proposed methodology versus the baseline using raw data.
@@ V-C Comparative Analysis With Other State-of-the-Art Models
Current Study 2025 Optimal Cascade Feature-Level Fusion 0.9936 0.9620 0.9584 0.9602
@@ VI Conclusion and Future Works
Autonomous vehicles have become a critical component of intelligent transportation systems. One of the most significant elements of autonomous vehicles is the CAN bus, which enables the connection of ECUs within the vehicle. Unfortunately, the CAN bus lacks built-in security mechanisms, making it vulnerable to various attacks. Existing solutions addressing this issue fail to focus on the key patterns of attacks and lack reliability and robustness. To address these shortcomings, this study proposes a model that integrates spatial and temporal information using cascade feature-level fusion optimized by a two-parameter genetic algorithm to tackle the dominant patterns of attacks. This approach ensures that zero-day attacks are detected by employing a method based on the intrinsic nature of the problem. The model demonstrates high reliability, with t-statistic and p-value of 20.51 and 5.09E-6, respectively. In addition, an AUC-ROC of 0.9783 emphasizes the reliability and robustness of the proposed model. Finally, the model is compared against various feature engineering and state-of-the-art models across two datasets. This comparison reveals that the proposed model outperforms other models and achieves outstanding accuracy and F1-score, establishing it as the best solution to this problem.

### v2

@@ I Introduction
  • Providing detailed results, performing statistical validation through paired t-tests to demonstrate the reliability and stability of the findings of the current study.
@@ II Related Works
TABLE I: Summary of Related Works
@@ III-A Datasets
TABLE II: Description of First Dataset [12]
@@
TABLE III: Description of Second Dataset [10]
@@ III-B Spatial Concept
TABLE VI: Feature Values Observed in CAN ID 0x490 During Spoofing Attack [12]
@@ III-C Temporal Concept
Figure 3: Time interval analysis for CAN ID 0x490 during attack [12].
@@ IV-B Spatial Concept
Raw Spatial Features 0.9681 0.0.7535 0.9076 0.8234
1D-CNN Spatial Features 0.9535 0.7149 0.7175 0.7162
Raw and 1D-CNN Spatial Features 0.9809 0.8766 0.8925 0.8845
xLSTM Spatial Features 0.9450 0.6606 0.6681 0.6643
Raw and xLSTM Spatial Features 0.9807 0.8758 0.8902 0.8830
Simple RNN Spatial Features 0.9398 0.6301 0.6395 0.6348
Raw and Simple RNN Spatial Features 0.9799 0.8726 0.8838 0.8782
GRU Spatial Features 0.9415 0.6405 0.6493 0.6449
Raw and GRU Spatial Features 0.9805 0.8763 0.8868 0.8815
LSTM Spatial Features 0.9420 0.6441 0.6513 0.6477
Raw and LSTM Spatial Features 0.9807 0.8765 0.8895 0.8830
@@
However, combining the module-generated features with raw spatial features leads to a significant enhancement in the F1-score by approximately 6% and in accuracy by around 1.5% across all models. This highlights not only the necessity of cascade feature-level fusion but also demonstrates its effectiveness. Additionally, cascade feature-level fusion enriched by xLSTM, LSTM, and 1D CNN, achieves accuracy higher than 0.9805, displaying the capability of these three models in this task.
@@ IV-C Temporal Concept
In the current subsection, the performance of raw temporal features, including Timestamp and CAN ID, as well as extracted temporal features, including Ratio and SE, and the combination of these four features are evaluated. Table IX shows that the model achieves an accuracy of approximately 94.64% when using only raw temporal features. In contrast, extracted temporal features alone result in a higher accuracy of 96.61%, indicating their effectiveness. However, combining extracted and raw features does not lead to a significant improvement, with the accuracy remaining around 94.66%. However, the combination of extracted temporal features with raw features does not show promising results, as expected. This is because the performance of the extracted features depends on the filter size, and there is a need to fine-tune this filter size. Therefore, the filter size is adjusted in the following subsection.
@@
Timestamp, CAN ID 0.9464 0.6817 0.6503 0.6656
Ratio, SE 0.9661 0.9198 0.6427 0.7566
Timestamp, CAN ID, 0.9466 0.6828 0.6524 0.6673
@@
9332 Timestamp, CAN ID, Data3, Data4, 0.9488
@@ IV-D Optimal Cascade Feature-Level Fusion
All Raw Features 0.9902 0.9335 0.9476 0.9405
1D-CNN-Based Optimal Cascade Feature-Level Fusion 0.9934 0.9604 0.9588 0.9596
xLSTM-Based Optimal Cascade Feature-Level Fusion 0.9934 0.9604 0.9592 0.9598
Simple RNN-Based Optimal Cascade Feature-Level Fusion 0.9934 0.9602 0.9588 0.9595
GRU-Based Optimal Cascade Feature-Level Fusion 0.9934 0.9609 0.9587 0.9598
LSTM-Based Optimal Cascade Feature-Level Fusion 0.9932 0.9588 0.9587 0.9588
@@ V-A Robustness and Reliability
This subsection examines the statistical evaluation of the model and analysis of its error rate, which are crucial for ensuring the reliability of the results. Table XIII presents the results of the 5x2cv paired t-test conducted to evaluate the statistical significance of the proposed model compared with the raw data. The t-statistic and p-value are equal to 20.51 and 5.09E-6, respectively. These results indicate that the proposed methodology is statistically significant.
TABLE XIII: Statistical Analysis of Proposed Optimal Cascade Feature-Level Fusion Method Using 5x2cv Paired T-Test
@@ V-C Comparative Analysis With Other State-of-the-Art Models
Current Study 2025 Optimal Cascade Feature-Level Fusion 0.9934 0.9604 0.9588 0.9596
@@ VI Conclusion and Future Works
Autonomous vehicles have become a critical component of intelligent transportation systems. One of the most significant elements of autonomous vehicles is the CAN bus, which enables the connection of ECUs within the vehicle. Unfortunately, the CAN bus lacks built-in security mechanisms, making it vulnerable to various attacks. Existing solutions addressing this issue fail to focus on the key patterns of attacks and lack reliability and robustness. To address these shortcomings, this study proposes a model that integrates spatial and temporal information using cascade feature-level fusion optimized by a two-parameter genetic algorithm to tackle the dominant patterns of attacks. This approach ensures that zero-day attacks are detected by employing a method based on the intrinsic nature of the problem. The model demonstrates high reliability, with t-statistic and p-value of 20.51 and 5.09E-6, respectively. Finally, the model is compared against various feature engineering and state-of-the-art models across two datasets. This comparison reveals that the proposed model outperforms other models and achieves outstanding accuracy and F1-score, establishing it as the best solution to this problem.
@@
VII Declaration of Generative AI and AI-Assisted Technologies in the Writing Process
During the writing process of this work, the authors used GPT-4o in order to enhance the text of the manuscript from a grammatical and lexical perspective. After using this tool/service, the authors reviewed and edited the content as needed and now take full responsibility for the content of the publication.